OpenStudy (janu16):

Christina has $103 earned from babysitting saved at home, and the amount is modeled by the function h(x) = 103. She reads about a bank that has savings accounts that accrue interest according to the function s(x) = (1.03)^(x − 1). After combining the two functions, the new function is g(x) = 103(1.03)^(x − 1). Using complete sentences, explain what this new function means.

OpenStudy (janu16):

@phi

OpenStudy (phi):

this looks peculiar h(x) = 103 h is one hundred three ??

OpenStudy (janu16):

yes

OpenStudy (phi):

ok. and this one s(x) = (1.03)^x − 1. is that \[ s(x) = 1.03^{x-1} \] or \[ s(x) = 1.03^x -1 \]

OpenStudy (janu16):

first one

OpenStudy (phi):

oh. then you should use parens around the (x-1)

OpenStudy (janu16):

okk

OpenStudy (phi):

I guess the combined function shows how much interest you get when you put in $103 into that bank. They don't say what x is, but I assume it measure time in months or years.

OpenStudy (janu16):

alright thanks so much.

OpenStudy (kinged):

It means that her deposit is getting 3% larger each period set by the bank (i.e. month, year...).